# Clustering-based Low-Rank Matrix Approximation: An Adaptive Theoretical Analysis with Application to Data Compression

### v1

@@ III-D The overall compression factor
3:$\beta_{\mathbf{C}_{k}}\leftarrow 0\quad\forall k\in\{1,\dots,K\}$
4:$N^{\text{eff}}\leftarrow 0$
5:for $k\leftarrow 1$ to $K$ do
[…]
8:         Increment $\mathbf{O}_{\mathbf{C}_{k}}[y:y+p,x:x+p]$ by $1$triangleright Mark overlapping pixels
9:     end for
@@
12:     $\beta_{\mathbf{C}_{k}}\leftarrow\frac{\text{overlapping\_pixels}_{k}}{\text{total\_pixels}_{k}}$triangleright Overlap proportion (Eq. 40)
13:     $N^{\text{eff}}_{\mathbf{C}_{k}}\leftarrow|\mathcal{P}_{\mathbf{C}_{k}}|(1-\beta_{\mathbf{C}_{k}})$triangleright Effective patches for cluster $\mathcal{C}_{k}$ (Eq. 41)
14:     $N^{\text{eff}}\leftarrow N^{\text{eff}}+N^{\text{eff}}_{\mathbf{C}_{k}}$triangleright Summing effective patches (Eq. 44)
@@
However, because $C$ is derived from independently determined local quantities, the precise equality can only occur under very special and highly regular conditions. That is, the local patches across the entire data matrix would need to be perfectly uniform, meaning that their characteristics such as texture, intensity, and structure are identical in every region. Second, the local compression parameters, including the number of patches $N_{\mathbf{C}_{k}}$, the effective patch counts $N^{\text{eff}}_{\mathbf{C}_{k}}$, and the singular value truncation ranks $r_{\mathbf{C}_{k}}$ for each cluster, must be exactly the same or follow a perfectly regular pattern. Third, the energy thresholds used in the local SVD must align exactly with the global energy distribution so that the overall energy preservation in the patch-based method yields an effective compression factor $CF_{\text{overall}}^{\text{eff}}(p,\alpha,K)$ that exactly matches one of the discrete values produced by $f(r_{g})$. Finally, the redundancy introduced by overlapping patches must be completely controlled so that the effective patch count $N^{\text{eff}}$ matches the theoretical prediction from the global formulation.
@@
where ||$\cdot$|| ⋅ | denotes the absolute value. $CF_{\text{threshold}}$ specifies the maximum allowable value of this relative difference so that the CFs are considered approximately equal. For example, if $CF_{\text{threshold}}$ is set to 0.15, this means that the CFs for global and cluster-based SVD must not differ by more than 15%.
@@
$CF^{\mathrm{eff}}_{\mathrm{overall}}(p,\alpha,K)=\frac{N^{\mathrm{eff}}(p)}{\displaystyle\sum_{k=1}^{K}r_{\mathbf{C}_{k}}(p)\,\phi_{\mathbf{C}_{k}}(p)},$
@@
$\phi_{\mathbf{C}_{k}}(p)=1+\frac{N_{\mathbf{C}_{k}}(p)+1}{p^{2}}.$
For each $k$ the function $N_{\mathbf{C}_{k}}(p)$ is non-increasing in $p$ while $p^{2}$ is strictly increasing; hence, each $\phi_{\mathbf{C}_{k}}(p)$ is strictly decreasing. Although the function $r_{\mathbf{C}_{k}}(p)$ (which is determined by Eq. 22) need not be monotonic in $p$, we assume that for every $k$, the combined function
$\phi_{\mathbf{C}_{k}}(p)=r_{\mathbf{C}_{k}}(p)\,\phi_{\mathbf{C}_{k}}(p)$
@@
$\phi_{\mathbf{C}_{k}}(p_{1})=r_{\mathbf{C}_{k}}(p_{1})\,\phi_{\mathbf{C}_{k}}(p_{1})\geq r_{\mathbf{C}_{k}}(p_{2})\,\phi_{\mathbf{C}_{k}}(p_{2})=\phi_{\mathbf{C}_{k}}(p_{2}).$
The strict decrease of $\phi_{\mathbf{C}_{k}}(p)$ due to the quadratic increase of $p^{2}$ and the decrease of $N_{\mathbf{C}_{k}}(p)$ ensures that even if $r_{\mathbf{C}_{k}}(p)$ does not decrease, the product $r_{\mathbf{C}_{k}}(p)\,\phi_{\mathbf{C}_{k}}(p)$ will tend to decrease as $p$ increases. Summing over $k$ yields
$\sum_{k=1}^{K}r_{\mathbf{C}_{k}}(p_{1})\,\phi_{\mathbf{C}_{k}}(p_{1})\geq\sum_{k=1}^{K}r_{\mathbf{C}_{k}}(p_{2})\,\phi_{\mathbf{C}_{k}}(p_{2}).$
Although $N^{\mathrm{eff}}(p)$ is non-increasing in $p$, we assume that the decrease in the denominator which is driven by the quadratic decay in $\phi_{\mathbf{C}_{k}}(p)$ dominates the decrease in $N^{\mathrm{eff}}(p)$. That is, for any $0<p_{1}<p_{2}$,
$\frac{N^{\mathrm{eff}}(p_{1})}{\displaystyle\sum_{k=1}^{K}r_{\mathbf{C}_{k}}(p_{1})\,\phi_{\mathbf{C}_{k}}(p_{1})}<\frac{N^{\mathrm{eff}}(p_{2})}{\displaystyle\sum_{k=1}^{K}r_{\mathbf{C}_{k}}(p_{2})\,\phi_{\mathbf{C}_{k}}(p_{2})}.$
@@
3:$N_{\text{eff}}\leftarrow 0$, $D\leftarrow 0$triangleright $D$ represents the denominator in (Eq. 43)
@@
5:     $N^{\text{eff}}_{\mathbf{C}_{k}}\leftarrow N_{\mathbf{C}_{k}}(1-\beta_{\mathbf{C}_{k}})$triangleright Effective patches for cluster $\mathcal{C}_{k}$ (Eq. 41)
6:     $N^{\text{eff}}\leftarrow N^{\text{eff}}+N^{\text{eff}}_{\mathbf{C}_{k}}$triangleright Summing effective patches (Eq. 44)
7:     $E_{\mathbf{C}_{k}}\leftarrow\sum_{i=1}^{\min\{N_{\mathbf{C}_{k}},p^{2}\}}\sigma_{i}^{2}$triangleright $E_{\mathbf{C}_{k}}$ is total energy for cluster $\mathcal{C}_{k}$
8:     $r_{\mathbf{C}_{k}}\leftarrow 0$, $P_{\mathbf{C}_{k}}\leftarrow 0$triangleright $P_{\mathbf{C}_{k}}$ is partial energy for cluster $\mathcal{C}_{k}$
@@
13:     $C_{\mathbf{C}_{k}}\leftarrow r_{\mathbf{C}_{k}}(N_{\mathbf{C}_{k}}+p^{2}+1)$triangleright $C_{\mathbf{C}_{k}}$ is cluster contribution for $k$
14:     $D\leftarrow D+C_{\mathbf{C}_{k}}$triangleright Summing cluster contributions for denominator
@@
16:$CF^{\text{eff}}_{\text{overall}}\leftarrow\displaystyle\frac{p^{2}N^{\text{eff}}}{D}$triangleright Effective overall $CF$ (Eq. 43)
@@
III-E Data matrix reconstruction
The compressed data matrix, $\widehat{\mathbf{A}}$, is reconstructed by combining the compressed patches from each cluster back into their original positions. In the compressed representation, each row of $\widehat{\mathbf{C}}_{k}$ corresponds to a vectorized patch extracted from the original matrix $\mathbf{A}$ at a specific position, determined by the top-left corner coordinates $(i_{k},j_{k})$ of the patch. For the $l^{\text{th}}$ patch in cluster $\mathcal{C}_{k}$, let $\Omega_{k,l}\subset[1,m]\times[1,n]$ be its spatial support, that is, the set of pixel coordinates covered by the patch. An indicator function is defined as
@@ III-E Data matrix reconstruction
Since the data matrix is split into overlapping patches, a pixel at location $(i,j)$ may be covered by several patches. Let $w_{ij}$ denote the number of patches covering pixel $(i,j)$. The final reconstructed image $\widehat{\mathbf{A}}\in\mathbb{R}^{m\times n}$ is computed as
@@
where $[\,\widehat{\mathbf{C}}_{k}\,]_{l}$ denotes the $l^{\text{th}}$ row of $\widehat{\mathbf{C}}_{k}$, the compressed representation of the $l^{\text{th}}$ patch in cluster $\mathcal{C}_{k}$, and $\Bigl{[}\operatorname{vec}^{-1}([\,\widehat{\mathbf{C}}_{k}\,]_{l})\Bigr{]}_{ij}$ is the pixel value at location $(i,j)$ in the reshaped $p\times p$ patch.
III-F Computation complexity
@@
IV Evaluation Metrics
This work uses several metrics to assess the quality of the reconstructed data matrix. The mean square error (MSE) quantifies the average squared difference between corresponding pixels of the original matrix $\mathbf{A}$ and its reconstruction $\widehat{\mathbf{A}}$ over a region of interest (ROI) $\Omega\subset[1,m]\times[1,n]$. It is computed as
@@ IV Evaluation Metrics
with a similar expression for the reconstructed data matrix. An EPI value of 1 indicates perfect preservation of edge information, while values closer to 0 suggest significant degradation.
V Experimental Setting
@@
Medical images are inherently large and contain complex structural details that must be preserved for accurate diagnosis, making efficient compression techniques crucial. Additionally, the high redundancy in medical imaging data makes it well-suited for low-rank approximation methods, which can reduce storage while maintaining critical diagnostic information. Therefore, we conducted experiments on medical images with different modalities to evaluate the effectiveness of global and cluster-based SVD in image compression and feature extraction. All computations were performed using a computer equipped with an 11th Gen Intel Core i7-1165G7 processor (2.80 GHz, 2.70 GHz base clock) and 16 GB RAM (15.6 GB usable). The system runs on a 64-bit operating system (x64 architecture). The global and cluster-based SVD were implemented using Python with libraries such as NumPy, SciPy, and OpenCV for matrix operations and image processing. To optimize computational efficiency, Joblib [54] was used for parallel computing, allowing k-means clustering and SVD calculations to be executed across multiple processors. The experimental results discussed in this section are specific to MRI. Outcomes for other imaging modalities such as ultrasound, CT scan, and X-ray are discussed in Appendix A. Readers are referred to Appendix A for detailed findings across these additional modalities.
V-A Dataset Description
@@
In this experiment we used the Brain Tumor Segmentation (BraTS) 2020 dataset, which is publicly available on Kaggle [55, 56, 57]. The BraTS 2020 dataset is widely recognized in medical image analysis, particularly for brain tumor segmentation and classification. It consists of high-resolution, multi-modal MRI scans—including T1-weighted, T1-contrast enhanced (T1ce), T2-weighted, and FLAIR sequences—acquired using standardized MRI protocols at multiple clinical centers. Each patient underwent a series of scans to capture a range of tissue contrasts, and the raw imaging data were preprocessed through steps such as skull stripping, co-registration, and intensity normalization to reduce inter-scanner variability and align the images to a common anatomical space. Expert radiologists then manually annotated the tumor regions, providing detailed segmentation masks that serve as the ground truth for evaluating image analysis algorithms. This dataset is ideal for evaluating compression and denoising techniques due to its high-resolution images and detailed structural features. We selected BraTS 2020 for its clinical relevance, as brain tumor detection is a critical task in medical imaging, and maintaining diagnostic accuracy during compression is essential. For this experiment, we used a single 2D slice from the middle of the MRI volume, ensuring consistency across different evaluations. The corresponding tumor segmentation mask was included in our analysis to assess the impact of compression separately on tumor and non-tumor region.
V-B Performance Evaluation
@@
The results presented in Figs. 6 to 8 were obtained using the MRI image shown in Fig. 5(a). To ensure consistency in the evaluation, all performance metrics and comparisons in this section were based on this specific image. The results for MRI image in Fig. 5(i) are included in Appendix A.
@@
Global SVD has a lower computational cost due to its simpler uniform compression strategy. However, as discussed, this approach is limited in its ability to preserve fine details, especially in the tumor region, and struggles to achieve higher compression without significant image degradation. On the other hand, cluster-based SVD incurs higher computational costs as the compression factor increases. This increase is due to the additional step of clustering and the independent SVD computations performed within each cluster as discussed in Section III-F. Despite these additional costs, cluster-based SVD outperforms global SVD in preserving structural details, features and maintaining superior image quality at different compression factors.
VI Discussion and Limitations

### v2

@@ 3.4 The overall compression factor
5:for $k\leftarrow 1$ to $K$ do
6:  $\mathbf{O}_{\mathbf{C}_{k}}\leftarrow\text{zeros}(\text{matrix\_shape})$ $\triangleright$ $\mathbf{O}_{\mathbf{C}_{k}}$ is the overlap map for cluster $\mathcal{C}_{k}$
8:   Increment $\mathbf{O}_{\mathbf{C}_{k}}[y:y+p,x:x+p]$ by 11 $\triangleright$ Mark overlapping pixels
9:  end for
10:  $\text{overlapping\_pixels}_{k}\leftarrow\sum(\mathbf{O}_{\mathbf{C}_{k}}>1)$ $\triangleright$ Count overlapping pixels
11:  $\text{total\_pixels}_{k}\leftarrow p^{2}|\mathcal{P}_{\mathbf{C}_{k}}|$ $\triangleright$ Total pixels in cluster $\mathcal{C}_{k}$ (note $|\mathcal{P}_{\mathbf{C}_{k}}|=N_{\mathbf{C}_{k}}$)
12:  $\beta_{\mathbf{C}_{k}}\leftarrow\frac{\text{overlapping\_pixels}_{k}}{\text{total\_pixels}_{k}}$ $\triangleright$ Overlap proportion (Eq. 40)
13:  $N^{\text{eff}}_{\mathbf{C}_{k}}\leftarrow|\mathcal{P}_{\mathbf{C}_{k}}|(1-\beta_{\mathbf{C}_{k}})$ $\triangleright$ Effective patches for cluster $\mathcal{C}_{k}$ (Eq. 41)
14:  $N^{\text{eff}}\leftarrow N^{\text{eff}}+N^{\text{eff}}_{\mathbf{C}_{k}}$ $\triangleright$ Summing effective patches (Eq. 44)
@@
However, because $C$ is derived from independently determined local quantities, the precise equality can only occur under very special and highly regular conditions. That is, the local patches across the entire medical image would need to be perfectly uniform, meaning that their characteristics such as texture, intensity, and structure are identical in every region. Second, the local compression parameters, including the number of patches $N_{\mathbf{C}_{k}}$, the effective patch counts $N^{\text{eff}}_{\mathbf{C}_{k}}$, and the singular value truncation ranks $r_{\mathbf{C}_{k}}$ for each cluster, must be exactly the same or follow a perfectly regular pattern. Third, the energy thresholds used in the local SVD must align exactly with the global energy distribution so that the overall energy preservation in the patch-based method yields an effective compression factor $CF_{\text{overall}}^{\text{eff}}(p,\alpha,K)$ that exactly matches one of the discrete values produced by $f(r_{g})$. Finally, the redundancy introduced by overlapping patches must be completely controlled so that the effective patch count $N^{\text{eff}}$ matches the theoretical prediction from the global formulation.
@@
where $|\cdot|$ denotes the absolute value. $CF_{\text{threshold}}$ specifies the maximum allowable value of this relative difference so that the CFs are considered approximately equal. For example, if $CF_{\text{threshold}}$ is set to 0.15, this means that the CFs for global and cluster-based SVD must not differ by more than 15%.
@@
$CF^{\mathrm{eff}}_{\mathrm{overall}}(p,\alpha,K)=\frac{N^{\mathrm{eff}}(p)p^{2}}{\displaystyle\sum_{k=1}^{K}r_{\mathbf{C}_{k}}(p)\,\psi_{\mathbf{C}_{k}}(p)},$
@@
$\psi_{\mathbf{C}_{k}}(p)=1+\frac{N_{\mathbf{C}_{k}}(p)+1}{p^{2}}.$
For each $k$ the function $N_{\mathbf{C}_{k}}(p)$ is non-increasing in $p$ while $p^{2}$ is strictly increasing; hence, each $\psi_{\mathbf{C}_{k}}(p)$ is strictly decreasing. Although the function $r_{\mathbf{C}_{k}}(p)$ (which is determined by Eq. 22) need not be monotonic in $p$, we assume that for every $k$, the combined function
$\phi_{\mathbf{C}_{k}}(p)=r_{\mathbf{C}_{k}}(p)\,\psi_{\mathbf{C}_{k}}(p)$
@@
$\phi_{\mathbf{C}_{k}}(p_{1})=r_{\mathbf{C}_{k}}(p_{1})\,\psi_{\mathbf{C}_{k}}(p_{1})\geq r_{\mathbf{C}_{k}}(p_{2})\,\psi_{\mathbf{C}_{k}}(p_{2})=\phi_{\mathbf{C}_{k}}(p_{2}).$
The strict decrease of $\psi_{\mathbf{C}_{k}}(p)$ due to the quadratic increase of $p^{2}$ and the decrease of $N_{\mathbf{C}_{k}}(p)$ ensures that even if $r_{\mathbf{C}_{k}}(p)$ does not decrease, the product $r_{\mathbf{C}_{k}}(p)\,\psi_{\mathbf{C}_{k}}(p)$ will tend to decrease as $p$ increases. Summing over $k$ yields
$\sum_{k=1}^{K}r_{\mathbf{C}_{k}}(p_{1})\,\psi_{\mathbf{C}_{k}}(p_{1})\geq\sum_{k=1}^{K}r_{\mathbf{C}_{k}}(p_{2})\,\psi_{\mathbf{C}_{k}}(p_{2}).$
Although $N^{\mathrm{eff}}(p)$ is non-increasing in $p$, we assume that the decrease in the denominator which is driven by the quadratic decay in $\psi_{\mathbf{C}_{k}}(p)$ dominates the decrease in $N^{\mathrm{eff}}(p)$. That is, for any $1<p_{1}<p_{2}$,
$\frac{N^{\mathrm{eff}}(p_{1})p^{2}_{1}}{\displaystyle\sum_{k=1}^{K}r_{\mathbf{C}_{k}}(p_{1})\,\psi_{\mathbf{C}_{k}}(p_{1})}<\frac{N^{\mathrm{eff}}(p_{2})p^{2}_{2}}{\displaystyle\sum_{k=1}^{K}r_{\mathbf{C}_{k}}(p_{2})\,\psi_{\mathbf{C}_{k}}(p_{2})}.$
@@
[…]
@@
5:  $N^{\text{eff}}_{\mathbf{C}_{k}}\leftarrow N_{\mathbf{C}_{k}}(1-\beta_{\mathbf{C}_{k}})$ $\triangleright$ Effective patches for cluster $\mathcal{C}_{k}$ (Eq. 41)
6:  $N^{\text{eff}}\leftarrow N^{\text{eff}}+N^{\text{eff}}_{\mathbf{C}_{k}}$ $\triangleright$ Summing effective patches (Eq. 44)
7:  $E_{\mathbf{C}_{k}}\leftarrow\sum_{i=1}^{\min\{N_{\mathbf{C}_{k}},p^{2}\}}\sigma_{i}^{2}$ $\triangleright$ $E_{\mathbf{C}_{k}}$ is total energy for cluster $\mathcal{C}_{k}$
8:  $r_{\mathbf{C}_{k}}\leftarrow 0$, $P_{\mathbf{C}_{k}}\leftarrow 0$ $\triangleright$ $P_{\mathbf{C}_{k}}$ is partial energy for cluster $\mathcal{C}_{k}$
@@
13:  $_{\mathbf{C}_{k}}\leftarrow r_{\mathbf{C}_{k}}(N_{\mathbf{C}_{k}}+p^{2}+1) \triangleright C𝐂$$_{\mathbf{C}_{k}}$ is cluster contribution for $k$
14:  $D\leftarrow D+C_{\mathbf{C}_{k}}$ $\triangleright$ Summing cluster contributions for denominator
@@
[…]
@@
3.5 Data matrix reconstruction
The compressed medical image, $\widehat{\mathbf{A}}$, is reconstructed by combining the compressed patches from each cluster back into their original positions. In the compressed representation, each row of $\widehat{\mathbf{C}}_{k}$ corresponds to a vectorized patch extracted from the original matrix $\mathbf{A}$ at a specific position, determined by the top-left corner coordinates $(i_{k},j_{k})$ of the patch. For the $l^{\text{th}}$ patch in cluster $\mathcal{C}_{k}$, let $\Omega_{k,l}\subset[1,m]\times[1,n]$ be its spatial support, that is, the set of pixel coordinates covered by the patch. An indicator function is defined as
@@ 3.5 Data matrix reconstruction
Since the medical image is split into overlapping patches, a pixel at location $(i,j)$ may be covered by several patches. Let $w_{ij}$ denote the number of patches covering pixel $(i,j)$. The final reconstructed image $\widehat{\mathbf{A}}\in\mathbb{R}^{m\times n}$ is computed as
@@
where $\widehat{\mathbf{C}}_{k}\,]_{l}$ denotes the $l^{\text{th}}$ row of $\widehat{\mathbf{C}}_{k}$, the compressed representation of the $l^{\text{th}}$ patch in cluster $\mathcal{C}_{k}$, and $\Bigl[\operatorname{vec}^{-1}([\,\widehat{\mathbf{C}}_{k}\,]_{l})\Bigr]_{ij}$ is the pixel value at location $(i,j)$ in the reshaped $p\times p$ patch.
3.6 Computation complexity
@@
4 Evaluation Metrics
This work uses several metrics to assess the quality of the reconstructed medical image. The mean square error (MSE) quantifies the average squared difference between corresponding pixels of the original matrix $\mathbf{A}$ and its reconstruction $\widehat{\mathbf{A}}$ over a region of interest (ROI) $\Omega\subset[1,m]\times[1,n]$. It is computed as
@@ 4 Evaluation Metrics
with a similar expression for the reconstructed medical image. An EPI value of 1 indicates perfect preservation of edge information, while values closer to 0 suggest significant degradation.
5 Experimental Setting
@@
Medical images are inherently large and contain complex structural details that must be preserved for accurate diagnosis, making efficient compression techniques crucial. Additionally, the high redundancy in medical imaging data makes it well-suited for low-rank approximation methods, which can reduce storage while maintaining critical diagnostic information. Therefore, we conducted experiments on medical images with different modalities to evaluate the effectiveness of global and cluster-based SVD in image compression and feature extraction. All computations were performed using a computer equipped with an 11th Gen Intel Core i7-1165G7 processor (2.80 GHz, 2.70 GHz base clock) and 16 GB RAM (15.6 GB usable). The system runs on a 64-bit operating system (x64 architecture). The global and cluster-based SVD were implemented using Python with libraries such as NumPy, SciPy, and OpenCV for matrix operations and image processing. To optimize computational efficiency, Joblib [34] was used for parallel computing, allowing k-means clustering and SVD calculations to be executed across multiple processors. The experimental results discussed in this section are specific to MRI. Outcomes for other imaging modalities such as ultrasound, CT scan, and X-ray are discussed in the appendix. Readers are referred to the appendix for detailed findings across these additional modalities.
5.1 Dataset Description
@@
In this experiment we used the Brain Tumor Segmentation (BraTS) 2020 dataset, which is publicly available on Kaggle [46, 5, 6]. The BraTS 2020 dataset is widely recognized in medical image analysis, particularly for brain tumor segmentation and classification. It consists of high-resolution, multi-modal MRI scans including T1-weighted, T1-contrast enhanced (T1ce), T2-weighted, and FLAIR sequences—acquired using standardized MRI protocols at multiple clinical centers. Each patient underwent a series of scans to capture a range of tissue contrasts, and the raw imaging data were preprocessed through steps such as skull stripping, co-registration, and intensity normalization to reduce inter-scanner variability and align the images to a common anatomical space. Expert radiologists then manually annotated the tumor regions, providing detailed segmentation masks that serve as the ground truth for evaluating image analysis algorithms. This dataset is ideal for evaluating compression and denoising techniques due to its high-resolution images and detailed structural features. We selected BraTS 2020 for its clinical relevance, as brain tumor detection is a critical task in medical imaging, and maintaining diagnostic accuracy during compression is essential. For this experiment, we used a single 2D slice from the middle of the MRI volume, ensuring consistency across different evaluations. The corresponding tumor segmentation mask was included in our analysis to assess the impact of compression separately on tumor and non-tumor region.
5.2 Performance Evaluation
@@
The results presented in Figs. 6 to 8 were obtained using the MRI image shown in Fig. 5(a). To ensure consistency in the evaluation, all performance metrics and comparisons in this section were based on this specific image. The results for MRI image in Fig. 5(i) are included in the appendix.
@@
Global SVD has a lower computational cost due to its simpler uniform compression strategy. However, as discussed, this approach is limited in its ability to preserve fine details, especially in the tumor region, and struggles to achieve higher compression without significant image degradation. On the other hand, cluster-based SVD incurs higher computational costs as the compression factor increases. This increase is due to the additional step of clustering and the independent SVD computations performed within each cluster as discussed in Section 3.6. Despite these additional costs, cluster-based SVD outperforms global SVD in preserving structural details, features and maintaining superior image quality at different compression factors.
6 Discussion and Limitations